A gas engine working on the Otto Cycle has cylinder bore of 200mm and stroke length of 250mm. The

clearance volume is 1570cm

3

. The pressure and temperature at the beginning of compression are 1 bar

and 27

0C respectively. The maximum temperature of the cycle is 1400

0C. Determine the pressure and

temperature at the salient points, the air standard efficiency, the work done and the mean effective

pressure. Also calculate the ideal power developed by the engine, if the number of working cycles per

minute is 500.


1
Expert's answer
2021-06-15T14:18:19-0400

"V_s=\frac{\pi d^2 h}{4}=0.00785~m^3,"

"r=\frac{V_c+V_s}{V_c}=6,"

"\eta=1-\frac{1}{r^{\gamma-1}}=51.16~\%,"

"T_2=T_1r^{\gamma-1}=614.3~K,"

"p_2=p_1r^{\gamma}=12.29~bar,"

"q_s=C_p(T_3-T_2)=1064.1~\frac{kJ}{kg},"

"w=\eta q_s=544.4\frac{kJ}{kg},"

"p_m=\frac{wm}{V_s-V_c}=\frac{wp_1V_s}{RT_1(V_s-V_c)}=2.73~bar."